Determine whether each of these functions is a bijection from R to R.

f (x) = (x2 + 1)/(x2 + 2)


1
Expert's answer
2021-09-10T12:09:46-0400

"f(x) = \frac{x^2+1}{x^2+2} \\


For \; x=-1 => f(x) = \frac{1+1}{1+2}= \frac{2}{3} \\


For\; x=1 => f(x) = \frac{1+1}{1+2}= \frac{2}{3}"

=> Not one-to-one function

=> Not a bijective (no inverse exists)